Write the correct name for each compound below.

  1. H3C-CH3:
  2. H3C-CH2-CH2-CH2-CH2-CH2-CH2-CH=CH2:
  3. H-C≡C-H:
  4. H-C-(CH3)3:
  5. H3C-CH=CH-CH3:
Answer #1

H3C-CH3 : ETHANE

H3C-CH2-CH2-CH2-CH2-CH2-CH2-CH=CH2 : NON-1-ENE

H-C\equivC-H : ETHYNE

H-C-(CH3)3 : (OR) CH3-CH(CH3)-CH3 : 2-METHYLPROPANE

H3C-CH=CH-CH3 : BUT-2-ENE

Note:

* Select longest possible C-C continuous chain as parent chain.

To indicate the number of carbons in the parent chain use root word.

No.of carbons in the parent chaine Root word
1 meth
2 eth
3 prop
4 but
5 pent
6 hex
7 hept
8 oct
9 non
10 dec

* For saturated compound use 'an' as primary suffix, unsaturated compound having double bond, use 'en' as primary suffix and for unsaturated compound having triple bond, use 'yn' as primary suffix.

* Use 'e' as secondary suffix for the compound having no functional group.

* Name the substitutents as prefix to the root word